Assume an economy with the following variables: YN = 24500; t = 18.5% ; YA = 21,000; G = 7250 ; Ta = 410. a) Compute the amount of taxes at natural GDP b) Compute the N.E.D (Natural Employment Deficit/Surplus). Express as a percentage of the natural GDP (YN). c) Policymakers decide to reduce/increase the size of the N.E.D to 2% of natural GDP. What will be the size of the N.E.D under these conditions? d) Given no change in the tax rate t, compute by how much they should cut G. e) Given no change in G, what should be the tax rate in order to accomplish the policy goal.
Answer #1

Question :

Answer :

Part A :

Taxes = 18.5% * 24500 = 0.185 * 24500 = 4532.5

Part B :

Natural Employment Deficit = Natural Real GDP - G

Natural Real GDP = t * Yn = 18.5% * 24500 = 4532.5

Natural Employment Deficit = 4532.5 - 7250 = - 2717.5

2717.5 = x% * 24500

X = 11.09%

Part C :

The new NED will be = 2% * 24500 = 490

New NED = - 490

Part D :

New NED = Natural Real GDP - G'

- 490 = 4532.5 - G'

G' = 4532.5 + 490 = 5022.5

Decrease in G = G - G' = 7250 - 5022.5 = 2227.5

If tax rate remains constant Government spending to be reduced by 2227.5

Part E : If G is constant, then

New NED = new t% *24500 - 7250

490 = 24500*t/100 - 7250

new t = 7250 + 490/245

new t = 31.59%

Increase in t = new t - t = 31.59 % - 18.5 % = 13.09%